OpenStudy (anonymous):

Solve for the variable in the equation 3.4 = -13.6 + (-3.4c) + 1.7c I don't know if i got it right, how does -1.7c look?

OpenStudy (anonymous):

I did something jacked up -_-

OpenStudy (amistre64):

3.4 = -13.6 + (-3.4c) + 1.7c ; clear the decimals by *10 34 = -136 -34c+17c ; combine like terms

OpenStudy (amistre64):

34 = -136 -34c+17c 136 136 ------------- 170 = -17c ; divide out the -17 -/17 -/17 ----------- -10 = c ; double chk of course

OpenStudy (amistre64):

3.4 = -13.6 -34 + 17 1 3.4 13.6 ----- 17 = -17 .... not quite lol id go with c = 10 then

OpenStudy (amistre64):

i got mixed up .... -10 is right, i just forgot how to multiply it back in :/

OpenStudy (anonymous):

hey sorry went afk had to help someone in school -_- about to go over give me a sec plz

OpenStudy (anonymous):

kk xD umm i think i got that lol